### v2

# Maximizing Battery Storage Profits via High-Frequency Intraday Trading

Abstract
Maximizing revenue for grid-scale battery energy storage systems in continuous intraday electricity markets requires strategies that are able to seize trading opportunities as soon as new information arrives. This paper introduces and evaluates an automated high-frequency trading strategy for battery energy storage systems trading on the intraday market for power while explicitly considering the dynamics of the limit order book, market rules, and technical parameters. The standard rolling intrinsic strategy is adapted for continuous intraday electricity markets and solved using a dynamic programming approximation that is two to three orders of magnitude faster than an exact mixed-integer linear programming solution. A detailed backtest over a full year of German order book data demonstrates that the proposed dynamic programming formulation does not reduce trading profits and enables the policy to react to every relevant order book update, enabling realistic rapid backtesting. Our results show the significant revenue potential of high-frequency trading: our policy earns $58$% more than when re-optimizing only once every hour and $14$% more than when re-optimizing once per minute, highlighting that profits critically depend on trading speed. Furthermore, we leverage the speed of our algorithm to train a parametric extension of the rolling intrinsic, increasing yearly revenue by $8.4$% out of sample.
keywords:
OR in energy , asset-backed trading , short-term power markets , dynamic programming , limit order book
journal: EJOR
[…]
[label1]organization=ETH Zurich, addressline=Rämistrasse 101, postcode=8092 Zurich, country=Switzerland
[…]
[label2]organization=Vrije Universiteit Amsterdam, addressline=De Boelelaan 1105, postcode=1081 HV Amsterdam, country=Netherlands
[…]
[label3]organization=University of Luxembourg, addressline=6 Rue Richard Coudenhove-Kalergi, postcode=L-1359, country=Luxembourg
[…]
[label4]organization=University of Bamberg, addressline=An der Weberei 5, postcode=96047 Bamberg, country=Germany
1 Introduction
Grid-scale battery energy storage systems (BESS) have emerged as a promising solution to deal with some of the short-term variability in renewable energy production. Batteries store excess energy generated during periods of high production and release it when generation is low, providing much-needed flexibility and stabilizing the grid (Martínez-Barbeito et al., 2023; Ullah et al., 2024). In addition, they can assist in managing peak demand, reducing the need for fossil fuel peaker plants (e.g., gas turbines) and enhance the overall efficiency of energy systems. The rapid increase in grid-scale storage installations underscores this potential (Saldarini et al., 2023; IRENA, 2017), with the projected adoption of BESS on the grid scale to increase significantly by 2030.
In liberalized electricity markets, the evaluation of viable business models for BESS is crucial to this transformation. The economic feasibility and revenue streams of these systems depend on their ability to participate in various short-term energy markets, engaging in energy arbitrage, frequency regulation, and demand response. Understanding the potential of these business models helps investors and legislators make the right decisions for a sustainable energy future and increases the adoption of storage technologies. In particular, informed trading decisions for intraday (ID) trading become increasingly relevant, with ID trading volumes growing each year (Koch and Hirth, 2019) and market participants shifting their trading towards short-term markets.
Finding good trading strategies requires extensive backtesting over long time horizons. However, accurately dealing with the fast pace of the continuous ID market, which operates at a millisecond frequency, while maintaining reasonable simulation times for long-term testing is challenging due to the resulting large number of required trading decisions.
In this paper, we propose and evaluate a fast automated intraday trading strategy that explicitly takes into account detailed order book dynamics, market rules, and technical limitations of the battery. As we argue below, the literature on the subject is rather scarce, as most authors adopt a rather stylized view of the intraday market avoiding most of its complexity and thereby yielding suboptimal strategies and a biased picture of revenue potentials.
Most relevant related literature on coordinated bidding in electricity spot markets focuses on the day-ahead market such as Fleten and Kristoffersen (2008) or treats the intraday market (IDM) as a single trading decision stage (Faria and Fleten, 2011; Löhndorf et al., 2013; Kongelf et al., 2019; Kraft et al., 2023). Applications range from pure trading strategies to asset-focused approaches, such as for hydro plant or battery system operation. Some publications consider a small number of repeated trading decisions on intraday markets (Aid et al., 2016; Löhndorf and Wozabal, 2023; Wozabal and Rameseder, 2020; Kuppelwieser and Wozabal, 2023).
Literature that exclusively focuses on (automated) trading strategies for the intraday market, while capturing its full complexity, is scarce. Some recent work presents approaches for single-, or multi-market trading strategies for storage. Jiang and Powell (2015) study short-term trading for a battery storage on real-time markets as a Markov Decision Process. Braeuer et al. (2019), and more recently, Seifert et al. (2024), present a European multi-market trading strategy for battery systems across three markets, trading at discrete time intervals and connecting the sequential reserve, day-ahead and intraday market decision stages. With a 4-hour time resolution, their ability to react to short-term changes in prices is very limited. Furthermore, the resulting smoothing of price spikes significantly reduces the profits a battery could extract from intraday markets.
The myopic rolling intrinsic (RI) trading strategy is nearly optimal for gas storage operation (Löhndorf and Wozabal, 2021). The strategy naturally translates to other storage types, like batteries. Consequently, the RI is frequently used as a benchmark for more sophisticated strategies (e.g., Bertrand and Papavasiliou, 2020; Boukas et al., 2021).
Recent work by Semmelmann et al. (2025) is most closely related to our work by presenting an evaluation of RI trading for a battery on the continuous intraday market. However, the authors do not take into account degradation cost in their optimization and solely rely on transaction data rather than the full order book and therefore miss critical information embedded in the full market data.
Generally speaking, there are only a handful of papers that model the intraday market in its full complexity and implement trading strategies that take into account the complete information in the order book and at the same time do not artificially discretize time. The few examples of such papers known to the authors include Bertrand and Papavasiliou (2020); Boukas et al. (2021); Kuppelwieser and Wozabal (2023). However, all of these papers are based on trading rules that are simple to execute. To the best of our knowledge, there is no single study of the rolling intrinsic policy or any other policy requiring a complex optimization for every single change in the limit order book.
This paper contributes to closing this gap by making the following main contributions:
  1. 1.
    We adapt the rolling intrinsic strategy to account for continuous trading on intraday markets for electricity, explicitly modeling every single order in the limit order book. The optimization is formulated as a mixed integer linear programming problem, taking into account all relevant aspects of the problem, including order placement costs, battery degradation, and detailed market rules.
  2. 2.
    We approximate the exact MILP formulation using a dynamic programming approach that is several orders of magnitude faster and nearly exact in our simulation settings. As a result, our work enables the first high-frequency algorithmic trading strategy for storage assets, enabling realistic trading at every relevant update of the order book. Furthermore, we provide a detailed comparison, in terms of speed and profit, of our method against the MILP solution.
  3. 3.
    In a numerical study, we conduct a detailed backtest of our strategy in the German intraday market over a full year using an order-by-order traversal of historical order book data. The results show the revenue potential of a BESS operating on the continuous intraday market.
    The computational speed of our proposed algorithm is essential for such a detailed analysis, and our findings demonstrate that the speed of trading is critical to maximize profits. Slower strategies, or generally less frequent trading, generate significantly lower profits by missing numerous trading opportunities due to their inability to make decisions at every relevant point in time. Our high-frequency strategy traverses the full year 2021 in around 86 minutes, solving the intrinsic optimization approximately 24 million times (4.6 solves per millisecond), submitting around 30k orders at the exchange. Furthermore, we evaluate the robustness of our strategy under optimization and battery parameter settings, providing a deeper understanding of the method’s performance and generalizability.
  4. 4.
    We parametrize the standard formulation of the RI, and show that our simulation speed can be used to easily find optimal parameters, generating additional profits of $8.5$, by slightly nudging the RI’s trading behavior at no additional risk.
  5. 5.
    Lastly, we publish an easy-to-use Python package to run RI simulations over extended periods of time, given a set of battery and dynamic programming parameters. This will allow other researchers or industry to realistically simulate high-frequency rolling intrinsic trading scenarios for a battery on the intraday market, thereby creating a strong benchmark which can be evaluated with minimal effort.
The remainder of this paper is organized as follows: Section 2 offers an overview of the European Power Market, setting the stage for our proposed trading strategy, detailed in Section 3. We present our results in Section 4, and discuss the details and implications of our findings. Finally, Section 5 concludes the paper and outlines directions for future research.
2 The European Power Market
The European electricity markets operate primarily within two major trading regions, each with their own dedicated exchange: the NordPool exchange, covering the Scandinavian region, and the EPEX exchange, which serves most of the rest of Europe.
The spot markets, which include day-ahead and intraday markets, are key components of the European Single Day-Ahead Coupling (SDAC) and Single Intraday Coupling (SIDC) initiatives. These initiatives have unified market operations across Europe, enabling bids submitted in one bidding zone to be seamlessly integrated into a shared order book and traded across other zones, provided there is sufficient cross-border transmission capacity. Price differences between bidding zones are thus primarily the result of constraints in interconnectors.
In this paper, we mainly focus on the German market, which is the largest electricity market in Europe and where trading is conducted mainly through the EPEX exchange. However, due to the ongoing convergence of the European electricity market designs, the proposed methods directly carries over to most other European markets.
The cascading range of future markets in Germany spans from long-term markets to day-ahead and intraday markets trading hourly and subhourly contracts. In continuous intraday trading, $60$, $30$ and $15$ minute products are trading up to $5$ minutes before delivery.
2.1 The Intraday Market for Power
The primary objective of short-term power market designs is to minimize system imbalances. The intraday (ID) market plays a crucial role as the final opportunity for market participants to adjust their positions in response to unforeseen changes in production and demand, often driven by updated weather forecasts. Participants use the ID market to fine-tune their positions from the day-ahead market, thereby minimizing deviations and reducing potential balancing costs charged by the transmission system operator (TSO). Significant price differences between the day-ahead and intraday markets create opportunities for participants to strategically hold back some positions from the day-ahead market and trade them on the intraday market for potential additional profits (Löhndorf and Wozabal, 2023; Seifert et al., 2024).
Intraday market designs in Europe vary by country, although there has been significant progress toward a more unified ID market structure through the Single Intraday Coupling (SIDC) initiative, which includes both continuous and recently also auction-based trading products (Epex Spot SE, 2024). In this paper we focus on the former.
In Germany, the intraday market opens in the afternoon following the clearing of the day-ahead market. Trading begins at 3 p.m. on the day before delivery, allowing a maximum of 32 hourly or 128 quarter-hourly products to be traded at any given time. For products listed in the shared SIDC order book, trading halts $60$ minutes before delivery, while trading for products solely in the German order book continues until $30$ minutes before delivery when it splits into the four German TSO regions, in each of which trading continues until $5$ minutes before delivery.
Limit orders submitted to the exchange are either immediately matched with corresponding buy or sell orders or, if unmatched, are stored in the limit order book (see Graf et al., 2024). A limit order is defined by its price, volume, validity, and order type (buy/sell). Figure 1 visualizes the process of a typical submission of a new limit order to the limit order book (LOB) of a product, i.e., delivery hour.
Figure 1: An exemplary state of the LOB at $T=1$ presented in the left panel. The right panel depicts the clearing of a new hypothetical buy order: A new buy order with a price of 50 €, which is higher than the lowest ask price, is added to the book at $T=2$. The quantity of this buy order (30 MW) is then cleared against the cheapest possible offers until either the whole order is fulfilled (as is the case in the figure) or until there are no offers with lower prices left. In this example, $22$ MW out of $30$ MW are cleared against the sell order with price 37 € and the remaining 30 MW $-\ 22$ MW $=8$ MW are cleared against the sell order with price 45 €. The remaining quantity of 12 MW $-\ 8$ MW $=4$ MW of the latter order stays in the order book. Note that the clearing is instantaneous, i.e., columns 2–4 in the right panel are purely illustrative and do not correspond to market states that can be observed by traders. Figure and caption adapted from Graf et al. (2024).
Continuous markets enable participants to continuously update and optimize their orders, allowing for real-time adjustments to evolving market conditions and reactions to forecast-updates. This flexibility makes them a crucial element in automated short-term trading strategies, where a rapid response to price changes is essential. In addition, continuous markets offer traders the opportunity to exploit short-lived market inefficiencies, enhancing their ability to capitalize on arbitrage opportunities throughout the trading day.
Figure 2: The binned distribution of time intervals between order submission and physical delivery for the German continuous intraday market in 2021. The teal bars illustrate the volume traded, while the red bars show the same data, counting the number of orders submitted. Both distributions exhibit an approximately exponential increase as the time to delivery shortens. Notably, there is a distinct gap between the 10-hour and 11-hour bins, where a significant increase in trading activity is observed.
A key challenge of trading in the continuous ID market is the lack of liquidity several hours before delivery. This results in wide bid-ask spreads and high price volatility. An example of this liquidity disparity is shown in Figure 2, which reveals that while a few hours before delivery typically thousands of orders are posted and cleared for a given product, this number drops to several hundred for products whose delivery is $10$ hours or longer away and to a handful of orders at the start of trading.
Low liquidity creates difficulties for automated trading strategies that aim to capitalize on price differences throughout the day. For instance, evening products traded during the morning hours often fail to reflect prevailing market prices accurately, as they are largely influenced by market makers rather than real supply and demand dynamics. This issue of liquidity is the main disadvantage of the continuous market, compared to its auction alternatives. It reduces the efficiency of price discovery and also increases trader’s risk as they may experience suboptimal pricing and reduced profitability when locking in positions early, as discussed more in-depth in e.g. (Graf et al., 2024).
Finally, Table 1 provides a concise summary of key statistics for the Central-Western European (CWE) Spot Market over recent years. The continuous market stands out as the dominant intraday trading platform. Specifically, the German continuous intraday market in 2021, which serves as the data source for this study, recorded 240.4 million order submissions (including order changes). This translates to an average of 7.6 orders placed per second. The sheer volume and frequency of transactions underscore the necessity of accurately modeling and responding to high-frequency market dynamics, to realistically simulate any trading strategy on this market. This calls for extremely fast and reactive trading. In the next section, we will propose an algorithm that can easily keep pace with the market’s speed.
2023 2022 2021 2020 2019 2018 2017
Day-Ahead (+16.9%) 419.3 (-8.6%) 358.8 (-4.5%) 392.7 (-8.4%) 411.1 (+3.1%) 448.8 (+5.3%) 435.2 413.2
ID-Auction (+9.0%) 9.41 (-1.9%) 8.63 (+0.0%) 8.8 (+20.5%) 8.8 (+7.4%) 7.3 (+30.8%) 6.8 5.2
ID-Continuous (+29.9%) 119.8 (+8.3%) 92.2 (+10.1%) 85.1 (+23.5%) 77.3 (+11.4%) 62.6 (+11.3%) 56.2 50.5
Table 1: Combined EPEX Spot trading volumes in TWh for the CWE region (AT, BE, DE/LU, FR, NL) taken from the EPEX Spot annual reports (EPEX SPOT SE, 2017-2023). The percentage-change to the respective last year is given in brackets. Clearly, intraday trading is gaining relevance compared to DA auctions, while the ID auction market is slowly picking up in activity.
3 An Efficient Rolling Intrinsic Policy for Continuous Markets
In this section, we will describe the intrinsic problem for continuous markets, discuss the rolling intrinsic policy as an improvement of the intrinsic strategy, and present an efficient implementation as a sequence of dynamic programming problems.
3.1 The Intrinsic Problem
The intrinsic strategy is a simple policy to maximize the rewards of a storage system, which is widely used in the literature on gas storage and in the industry to optimize storage assets (e.g. Gray and Khandelwal, 2004b, a; Lai et al., 2010, 2011; Löhndorf and Wozabal, 2021). In the extant literature, the intrinsic problem is set in a perfectly liquid futures market, trading $T\in\mathds{N}$ futures of a commodity that can be stored and which trade for prices $P_{t}$. In order not to overcomplicate notation, we assume here that future contracts are for delivery in consecutive time periods $[0,1),[1,2),\dots,[T-1,T)$ and that there are no overlapping contracts. To fix ideas, one can imagine a battery storage system trading futures for delivery in the $24$ hours of the next day.
The storage operator decides about traded quantities $q_{t}$ of these contracts such that volumes bought are transferred into the storage, while sold volumes are physically fulfilled from the storage. In particular, in the simplest case, the intrinsic optimization boils down to the following linear problem of optimizing the storage level $(s_{t})_{t=1}^{T}$ and the future positions $(f_{t})_{t=1}^{T}$ in the following way
$\begin{array}[]{lll}\underset{q_{t}}{\max}&\sum\limits_{t=1}^{T}P_{t}(-q_{t})\\ \text{s.t.}&f_{t}=f^{0}_{t}+q_{t},&\forall t=1,\dots,T\\ &f_{t}\in[\underline{f},\bar{f}],&\forall t=1,\dots,T\\ &s_{t}=s_{t-1}+f_{t},&\forall t=1,\dots,T\\ &s_{t}\in[0,\bar{s}],&\forall t=1,\dots,T,\end{array}$ (6)
where $s_{0}$ and $(f_{t}^{0})_{t=1}^{T}$ are the initial storage level and the initial future positions, respectively, which are data to the problem. In the above problem, positive $q_{t}$ represent buying decisions while negative ones represent sells. Furthermore, $\bar{s}$, $\underline{f}$, and $\bar{f}$ are physical bounds for the maximum storage level (MWh) as well as for withdrawal and injection (MW). The respective constraints ensure that the position is physical in the sense that traded quantities can be accommodated in the storage without violation of the limits on injection and withdrawal.
The above linear problem is the simplest form of the intrinsic, which assumes a perfectly efficient storage, liquid markets, and no cost for operating the storage. Furthermore, the problem is myopic in the sense that the decision maker uses all the flexibility of the storage instantly without anticipating future changes in prices, which would make it more advantageous to wait before trading immediately. The myopic nature of the intrinsic policy makes it suboptimal and motivates its name originating from the intrinsic value of a financial option contract, i.e., the value an option has when it is immediately exercised, not taking into account the value of waiting (the extrinsic value). Problem (6) is the version of the intrinsic problem prevailing in the literature.
Next we will extend (6) to a policy for a BESS trading on continuous intraday power markets, explicitly taking into account order book information, trading costs, and a simple linear model for efficiency and degradation losses of batteries. In order to do so, we introduce a set of future contracts $\mathcal{T}_{t^{*}}$ at time $t^{*}$, for delivery of electricity in non-overlapping time periods in the future. The time when the earliest contract goes into delivery is denoted by $t_{0}$. Note that the formulation could easily be extended to future contracts that overlap, such as quarter-hourly and hourly contracts. For simplicity, in the following we only focus on the intrinsic at a singular time $t^{*}$, denoting $\mathcal{T}_{t^{*}}$ simply as $\mathcal{T}$.
For each contract $t\in\mathcal{T}$, there is an order book $\mathcal{O}_{t}=\mathcal{O}_{t}^{+}\cup\mathcal{O}_{t}^{-}$ with $\mathcal{O}_{t}^{+}$ containing all the currently active asks and $\mathcal{O}_{t}^{-}$ containing all the currently active bids for contract $t$. For each of these orders $i\in\mathcal{O}_{t}$, we define a limit price $P_{i}$ and a quantity $Q_{i}>0$. Furthermore, we introduce efficiency factors $\eta^{+}$, $\eta^{-}\in(0,1]$ for charging and discharging, respectively, as well as a volume-based trading cost $\nu_{\text{trade}}$ for matching orders on the market.
Batteries have a limited life. Roughly speaking, there are two effects causing battery degradation: calendar age, i.e., the loss of capacity with time even without active use of the battery, and cycling, i.e., the wear and tear caused by injection and withdrawal of energy. Since we are looking at short-term planning, we focus on the latter and model a linear cycling (i.e. degradation) cost $\nu_{\text{deg}}$ in €/MWh of withdrawn energy. The intuition behind this choice is that cells will have to be replaced after a certain number of cycles, which induces a replacement cost. With this information, we calculate a proportional per cycle and ultimately a per MWh cost for using the battery. We note that actual degradation depends not only on the amount of cycled energy but also on the depth of discharge, temperature, and several other factors (see e.g. Xu et al., 2017; Gräf et al., 2022). However, for the sake of simplicity, we focus on the amount of cycled energy which is the most important factor. Combining the cost of trading and the cost of degradation, we define $\nu=\nu_{\text{trade}}+\nu_{\text{deg}}$ as the combined variable cost in units of €/MWh.
Until this point, all discussed extensions of (6) can still be modeled as a linear program (LP). However, there are two aspects of intraday trading that introduce non-linearities. Firstly, trade quantities on electricity exchanges are not continuous variables but multiples of the minimum trading unit $u$, which implies that they are of the form $q_{i}=k_{i}u$ with $k_{i}\in\mathds{N}_{0}$. Secondly, efficiencies smaller than $1$ coupled with negative prices make it profitable to spend energy by simultaneously charging and discharging. More specifically, for a single product $t$ and orders $i\in\mathcal{O}^{+}_{t}$, $j\in\mathcal{O}^{-}_{t}$, buying a quantity $q$ using order $i$ and immediately re-selling the power $\eta^{-}(\eta^{+}q)$ that remains after accounting for efficiency losses using order $j$ yields a positive profit if
$-qP_{i}+\eta^{-}(\eta^{+}q)P_{j}>0\Leftrightarrow(\eta^{-}\eta^{+})P_{j}>P_{i}.$
Since $P_{i}>P_{j}$ by definition, this only can happen if both prices are negative. However, because it is physically not possible to charge and discharge the battery at the same time, such trades have to be prevented.
With these preparations, we can write the modified intrinsic as the following mixed integer linear problem (MILP):
$\underset{f_{t},s_{t},\alpha_{t},q_{i},k_{i}}{\max}\quad$ $\sum\limits_{t\in\mathcal{T}}\left(\sum\limits_{i\in\mathcal{O}_{t}^{-}}(P_{i}-\nu)q_{i}-\sum\limits_{i\in\mathcal{O}_{t}^{+}}(P_{i}+\nu)q_{i}\right)$ (7)
s.t. $0\leq q_{i}\leq Q_{i},$ $\forall i\in\mathcal{O}_{t}$
$q_{i}=k_{i}u,$ $\forall i\in\mathcal{O}_{t}$
$f_{t}^{+}=\sum_{i\in\mathcal{O}_{t}^{+}}q_{i},$ $\forall i\in\mathcal{O}_{t}^{+}$
$f_{t}^{-}=\sum_{i\in\mathcal{O}_{t}^{-}}q_{i},$ $\forall i\in\mathcal{O}_{t}^{-}$
$f_{t}=f^{0}_{t}+f_{t}^{+}-f_{t}^{-},\qquad$ $\forall t\in\mathcal{T}$
$f_{t}\in[\underline{f},\bar{f}],$ $\forall t\in\mathcal{T}$
$f_{t}=i_{t}-w_{t},$ $\forall t\in\mathcal{T}$
$i_{t}\in[0,\alpha_{t}\bar{f}],$ $\forall t\in\mathcal{T}$
$w_{t}\in[0,-(1-\alpha_{t})\underline{f}],$ $\forall t\in\mathcal{T}$
$s_{t}=s_{t-1}+\eta^{+}i_{t}-\frac{1}{\eta^{-}}w_{t},$ $\forall t\in\mathcal{T}$
$s_{t}\in[0,\bar{s}],$ $\forall t\in\mathcal{T}$
$\alpha_{t}\in\{0,1\},\;k_{t}\in\mathds{N},\;\nu\in\mathbb{R}_{\geq 0}$ $\forall t\in\mathcal{T}$
where the $\alpha_{t}$ are binary variables which are equal to $1$ in case the decision for contract $t$ is to buy and 00 if the decision is to sell.
Note that, due to the above discussion, strictly speaking, the binary variables $\alpha_{t}$ are only required for hours $t$ where both $\mathcal{O}_{t}^{+}$ and $\mathcal{O}_{t}^{-}$ contain orders with negative prices. Since $u$ is usually rather small and negative prices do not occur too often, typically linear relaxations of the above problems are good and yield near-optimal solutions in the root node of the branch-and-bound trees, making the problem a comparably easy MILP in most realistic cases.
3.2 The Rolling Intrinsic
Input: Initial storage $s_{0}$, start-time $t_{start}$ and stop-time of simulation $t_{end}$, limit order book information $\mathcal{T}_{t^{*}}\ \forall t^{*}\in[t_{start},t_{end}]$, solve-frequency of intrinsic, other parameters $(\bar{f},\underline{f},\nu,\dots)$
Output: Final battery schedule during simulation runtime, cumulated trade profits
for time instant $t^{*}\in[t_{start},t_{end}]$ do
     Solve the intrinsic according to equation (7)
     Update all positions $f_{t},\forall t\in\mathcal{T}_{t^{*}}$
     Log profits/losses from trading
end for
Algorithm 1 The rolling intrinsic
The rolling intrinsic policy (RI), as outlined in Algorithm 1 was originally introduced in Gray and Khandelwal (2004b) and is a dynamic extension of the intrinsic value discussed in the last section. Starting from an initial storage state $s_{0}$ and market positions $(f_{t})_{t\in\mathcal{T}}$ acquired in previous periods, the RI repeatedly checks for chances of profitable rebalancing by re-running the intrinsic policy. Although the resulting decisions are still myopic, the RI clearly represents an improvement over the static intrinsic policy that does not adapt positions at all. Furthermore, the myopic nature of the decisions has the advantage that the RI does not speculate, but only enters immediately profitable positions and therefore does not run the risk of accumulating losses. For these reasons and because of its conceptual simplicity and relatively low computational cost, the RI has gained widespread industry adoption.
3.3 A Dynamic Programming Formulation of the Rolling Intrinsic Policy
In this section, we describe how to reformulate problem (7) to a dynamic programming formulation that can be solved by an efficient implementation of the dynamic programming algorithm. The aim of the reformulation is to make faster decisions to keep up with the pace of the intraday market, where typically many orders arrive per second, requiring constant rapid reevaluation of one’s position.
The main idea of the reformulation is to assign each product $t\in\mathcal{T}$ a stage in which the decisions to accept orders $i\in\mathcal{O}_{t}$ are taken while the storage level is the state variable that connects the stages. Hence, in the language of dynamic programming, the decisions $q_{i}$, $k_{i}$, $\alpha_{t}$, and $f_{t}$ are the actions, and the resulting storage state $s_{t}$ is the argument of the value function of the next stage.
We assume without loss of generality that $\mathcal{T}_{t}=\{1,\dots,T\}$ set $V_{T+1}\equiv 0$ and for $t=1,\dots,T$ define the value function as
[…]
This formulation might not seem very natural, since in reality the decisions in all stages of the above problem happen simultaneously instead of sequentially as usually is the case in dynamic programs. However, this perspective allows for a faster solution than solving the monolithic MILP or LP as we will argue below.
We propose solving the problem by an application of the dynamic programming algorithm, discretizing the actions $f_{t}$ into steps $v\cdot u$ ($v\in\mathds{N}$) apart, where $u$ is the minimal tradable quantity. Furthermore, we introduce a function $\pi_{t}$ that encodes the current order book information and returns the cost or revenue of buying or selling on the intraday market, taking into account trading costs. Furthermore, we define a state transition function
$S(s_{t-1},f_{t})=\begin{cases}s_{t-1}+\eta^{+}f_{t},&\text{if }f_{t}>0\\ s_{t-1}+\frac{f_{t}}{\eta^{-}},&\text{otherwise.}\end{cases}$
With these preparations, we can now re-write problem (20) as
[…]
where the bounds in the first constraint $k_{t}$ enforce the storage energy limits, while the second interval enforces the power limits. Given $V_{t+1}$, solving (21) boils down to evaluating the objective for all feasible $k_{t}$, i.e., to a small number of arithmetic operations.
Note that the above reformulation of (20) to (21) automatically takes care of the issue of simultaneously accepting buy and sell orders in the case of negative prices, and, by fixing the trading step-size $u$ to a multiple of the minimum bid-size, naturally yields feasible position sizes.
To solve (21), we need to know the value functions $V_{t+1}$. In line with the usual dynamic programming algorithm, we calculate $V_{t}$ going backward in time: $V_{T+1}$ is the known boundary condition, which makes it possible to calculate $V_{T}$ in any given state. However, due to efficiencies $\eta^{\pm}<1$, the trades that live on a finite grid translate to storage levels that do not sit on a regular grid, especially after several rounds of injections and withdrawals.
Therefore, unlike classic applications of the dynamic programming algorithm, we do not define $V_{T}$ to have a finite domain, but instead on the whole interval of possible storage states $[0,\bar{s}]$. In order to obtain the value function, we discretize the possible storage states $[0,\bar{s}]$ to a finite grid $G\subset[0,\bar{s}]$ and $\lvert G\rvert=m$. We then approximate the value function $V_{T}$ by evaluating it on the grid and linearly interpolating for points $s\notin G$.
In particular, we denote the resulting approximations by $\tilde{V}_{t+1}$ and describe our linear approximation between the neighboring discretized gridpoints $s_{i},s_{i+1}\in G$ as follows:
$\tilde{V}_{t+1}(s_{t})=\frac{m(s_{i+1}-s_{t})}{\bar{s}}V_{t+1}(s_{i})+\frac{m(s_{t}-s_{i})}{\bar{s}}V_{t+1}(s_{i+1}),$ (22)
for $s_{i}\leq s_{t}\leq s_{i+1}$. More sophisticated interpolations, taking into account, for example, the curvature of the value function (given by the buy and sell price stacks), could potentially enhance the precision of our optimization. However, since such approaches would be computationally more costly and the obvious choices such as quadratic interpolations or cubic splines did not show any improvements in our numerical experiments, we remain with the linear approximation above.
Once we have obtained $\tilde{V}_{T}$, we can solve the problem that defines $V_{T-1}$ with $V_{T}$ replaced by $\tilde{V}_{T}$. We then repeat the process until we have obtained the approximations $\tilde{V}_{t}$ of all the value functions. Note that, in principle, $G$ can be chosen independently for every $t$ and can also change between various solutions of the intrinsic problem. For the sake of simplicity, we keep the grid constant, i.e. $G_{t^{*},t}\equiv G$.
Algorithm 2 displays an overview of one such dynamic program solution run of the intrinsic problem at time $t^{*}$, where $\mathcal{X}_{t}$ is the set of all possible actions $k_{t}$ as defined in equation (21).
Algorithm 2 DP at runtime $t^{*}$
Input: $f_{t}^{0}$, $\mathcal{O}_{t}$ for all tradable products $t\in\mathcal{T}$, other parameters $(v,s_{0},\nu,G,\dots)$
Output: Updated market positions $f_{t}$
Set $V_{T+1}\equiv 0$.
Backwards Pass
for $t=T,\dots,t_{0}$ do
     for each state $s\in G$ do
         $V_{t}(s)\leftarrow\max_{k_{t}\in\mathcal{X}_{t}(s,f_{t}^{0})}\left\{\pi_{t}(k_{t}u)+\tilde{V}_{t+1}(S(s,k_{t}u))\right\}$
     end for
end for
Forward Pass
for $t=t_{0},\dots,T$ do
     $\hat{f_{t}}\leftarrow\operatorname*{arg\,max}_{k_{t}\in\mathcal{X}_{t}(s,f_{t}^{0})}\left\{\pi(k_{t}u)+\tilde{V}_{t+1}(S(s_{t-1},k_{t}u))\right\}$
     $f_{t}^{0}\leftarrow\hat{f_{t}}$
     $s_{t}\leftarrow s_{t-1}+\hat{f_{t}}$
end for
The approximation of the value functions $\tilde{V}_{t}$ introduces an error relative to the exact MILP formulation of the problem. In the following proposition, we show that the domain of the true functions $V_{t}$ is not actually continuous and that consequently there is a grid $G$, which makes the approximation and thereby the DP formulation of the problem exact.
Proposition 1.
For all $t\in\left\{1,\dots,T\right\}$, there is a finite grid $G_{t}$ that contains all possible storage states at the beginning of period $t$. Choosing $G$ to approximate the value functions alleviates the need for interpolation and makes the approximation exact.
Proof.
Suppose that the initial state of charge (SoC) is $s_{0}$, we set $v=1$ and look at the simplest case of $f_{t}^{0}=0$ for all $t\in\mathcal{T}$. The possible storage states at the beginning of stage $t_{0}+1$ are
$G_{t_{0}+1}=\left\{s_{0}+uk_{1}^{+}\eta^{+}-\frac{uk_{1}^{-}}{\eta^{-}}:k_{1}^{\pm}\in\mathcal{K}_{1}^{\pm}\right\}$
and $\mathcal{K}_{1}^{+}=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\min\left(\frac{\bar{s}-s_{0}}{\eta^{+}},\bar{f}\right)\right\}$ and $\mathcal{K}_{1}^{-}=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\min\left(s_{0}\eta^{-},-\underline{f}\right)\right\}$.
Defining the sets $G_{t}$ and $\mathcal{K}_{t}^{\pm}$ going forward in time, for period $t>t_{0}$ we thus have, given that $\mathcal{K}_{t-1}^{\pm}$ is already known,
$G_{t}=\left\{s_{t}=s_{0}+uk_{t}^{+}\eta^{+}-\frac{uk_{t}^{-}}{\eta^{-}}:k_{t}^{\pm}\in\mathcal{K}_{t}^{\pm},0\leq s_{t}\leq\bar{s}\right\}$
with
$\mathcal{K}_{t}^{+}$ =$=\left\{k\in\mathds{N}_{0}:k\leq\frac{(\bar{s}-s_{0})+u\max(\mathcal{K}_{t-1}^{-})(\eta^{-})^{-1}}{u\eta^{+}}\right\}$
$\mathcal{K}_{t}^{-}$ =$=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\eta^{-}\left(s_{0}+\eta^{+}u\max(\mathcal{K}_{t-1}^{+})\right)\right\}.$
Note that the way $G_{t}$ and $\mathcal{K}_{t}^{\pm}$ are defined takes care of repeated charging and discharging in periods $t_{0},\dots,t$ and therefore covers all possible storage states. ∎
Choosing the finite grids described above would ensure that the DP solution is exact and equivalent to the MILP formulation, as the value function only needs to be evaluated at points in the grid and there is no need for approximation. This comes at the cost of computational efficiency, as choosing $G$ to be exact, would yield a very fine grid for cases where there are many tradable products $\mathcal{T}$ and $\eta^{\pm}<1$.
4 Results
In this section, we present the results of a numerical case study. In Section 4.1, we discuss the setting and implementation of the case study. In Section 4.2, measure the precision of our DP solution and compare the runtimes with the exact MILP formulation, arguing that the DP leads to runtimes that are orders of magnitude faster. In Section 4.3, we present the results of our policy for a whole year of trading and measure the impact of trading speed on performance, while in Section 4.4, we examine how changes in storage and optimization parameters affect the results.
4.1 Setting and Mode of Comparison
Unless otherwise stated, for our case study, we use a BESS with an energy-to-power ratio (duration) of 1 hour, setting $\bar{f}=-\underline{f}=10$ MW and $\bar{s}=10$ MWh. We assume efficiency losses of $\eta^{+}=\eta^{-}=0.95$, i.e., a round trip efficiency of $\eta^{+}\eta^{-}\approx 0.9$ similar to Cole and Karmakar (2023); Regelleistung Online (2024).
We use linear degradation costs of $\nu_{\text{deg}}=4$/MWh, i.e., a round-trip degradation cost of $8$/MWh for every MWh of electricity sold by the battery. We motivate this value by Goldman Sachs (2023) who project the cost of battery replacement in 2030 to be around $60$k €/ MWh, a 10-year battery warranty, and approximately $2$ charging cycles per day. Dividing the replacement costs by cycles gives the cost of one full cycle, which allows us to infer the induced per MWh degradation cost.
Furthermore, we choose $\nu_{\text{trade}}=0.09$/MWh, reflecting the current EPEX volume-based trading fees per each matched order. In total, the variable cost $\nu=\nu_{\text{trade}}+\nu_{\text{deg}}$ thus takes the value of $\nu=4.09$/MWh.
All results are generated using historical 2021 order book data from the EPEX Spot continuous intraday market for Germany, focusing on hourly products. The exchange allows for price increments of $0.01$ and volume increments at $0.1$ MWh, defining the minimum action size on the market as $u=0.1$ MWh. The discretization of the dynamic programming action is defined at the most accurate minimum value of $v=1$ throughout.
Clearly, in real-life trading, there will always be some delay between two consecutive optimizations. We explicitly model this delay $\Delta t$ and conceptually split it into two components: the solve-time and the technical delay. The solve time represents the time required to find an optimal solution for a given intrinsic problem. The technical delay accounts for various operational latencies, such as communication delays with the exchange, delays caused by internal database and trading systems, and other post-solve delays before the submitted orders are matched on the exchange.
$t^{*}$$t^{*}+\Delta t$Solve TimeTechnical DelayWaiting Duration
If the combined delay exceeds $1$ ms, this has two effects on trading: Firstly, there is the risk that an order placed by our policy may not be matched since its intended counterpart either expired or was cleared against other market participants’ orders in the meantime. We deal with this by submitting all-or-none orders that expire if they are not immediately cleared, which might lead to non-physical positions, i.e., market positions that cannot be matched by a feasible schedule of injections and withdrawals. In such a situation, further trades are required to correct the position so that the final battery schedule becomes physically feasible. However, in the way the intrinsic problem is set up, such positions would be corrected in the next call of the intrinsic problem, forcing the RI to return to a physical position as quickly as possible.
The second issue arising from delays is that the policy might not be able to react to every single order book update. In particular, we assume that the next intrinsic is solved only at $t^{*}+\Delta t$ observing the updated order book state at this time. Hence, if there are updates to the order book in the time interval $(t^{*},t^{*}+\Delta t)$, the policy will not be able to react immediately to it.
Finally, we remark that when backtesting, we correctly account for changes in the order book caused by our actions. This in particular means that orders cleared against the orders of the policy are available shorter than in the actual history, and consequently that subsequent clearing of orders by other markets participants changes.
Although we take great care to model the market correctly, a limitation of the backtesting experiments is that, by the very nature of our analysis and the available data, we cannot take into account the effect that the orders placed by the strategy would have had on the behavior of other market participants. However, this drawback is inherent in the idea of backtesting and cannot be easily corrected.
All simulations were implemented in C++ and run on a single AMD EPYC 9654 CPU, with the MILP solved using Gurobi v11.0.3 for C++ (Gurobi Optimization, LLC, 2024). Due to the simple nature of each single intrinsic optimization, solving the MILP on multiple threads does not lead to faster simulation times. We publish or code in the Python package BitePy, which allows users to easily run our DP simulations over a Python interface. See A for details.
4.2 Comparison between MILP and DP Rolling Intrinsic
To evaluate the effectiveness of solving the intrinsic rolling problem using the DP approach instead of the MILP approach, we start by comparing the two methods. The MILP consistently provides the exact solution to the optimization problem (7), while the DP solution generally slightly diverges. Note that, due to the RI’s myopic nature and the path dependence of differing solutions, following the exact solution does not necessarily guarantee larger profits. For this reason, although inexact in solving single intrinsic problems, the DP-based rolling intrinsic strategy can be more profitable than the MILP formulation, as evidenced by the February and April results in Table 2.
We choose the most clinical setting for our comparison, by artificially fixing the solve time and technical delay to 00 ms for both approaches. This results in a policy that ignores technical delays and solves the intrinsic optimization for each relevant order book update in such a way that the resulting all-or-none limit orders are matched instantaneously with the exchange after each solve.
Table 2 reports this comparison for the first week of February, April, July, and October 2021. All three DP methods use a constant equidistant grid $G$, with storage discretizations $m=101$, $m=51$ and $m=11$. In summary, we can say that even without delays, the DP achieves comparable rewards to MILP and significantly outperforms it in terms of runtime by a factor of 100-1000 for the coarser storage discretization $m=11$. While the DP simulation times remain constant, MILP simulation times vary, even though the number of intrinsic solves is almost equal, as the complexity of the MILP problem is highly influenced by current market conditions. Interestingly, a finer storage discretization does not provide a clear advantage over coarser discretizations in this experiment, but shows an increase in profits for our yearly simulation results, as evidenced by the following sections.
Reward Sim-Time Cycles/Day Solves Traded Vol.
[€] [h] [MWh]
Feb 01-14 MILP 7895 6.77 2.3 959900 1901
DP-101 7538 0.43 2.3 959900 1825
DP-51 8051 0.22 2.2 959800 1905
DP-11 7851 0.05 2.2 959500 1878
Apr 01-14 MILP 9487 15.81 2.5 733000 1856
DP-101 9434 0.34 2.5 733100 1817
DP-51 9353 0.17 2.5 733000 1792
DP-11 9543 0.04 2.5 733200 1872
Jul 01-14 MILP 11565 17.07 2 901000 1948
DP-101 11161 0.43 2 901000 1873
DP-51 11158 0.22 2 900900 1884
DP-11 11181 0.05 2 901000 1899
Oct 01-14 MILP 28553 56.93 2.6 1012900 3350
DP-101 28134 0.47 2.6 1012500 3231
DP-51 27952 0.24 2.5 1012600 3190
DP-11 28014 0.06 2.5 1012600 3236
Table 2: Performance comparison for four representative weeks in 2021 between the rolling intrinsic solved using the benchmark MILP and three DP approaches with different approximations due to their value function granularities. All results are rounded to significance.
4.3 Yearly High-Frequency Rolling Intrinsic Trading
Figure 3: The cumulative reward of the intrinsic rolling over the entire year 2021 for various intrinsic-solve frequencies. The inset shows the battery operation over the span of two arbitrary selected days for the strategy which solves the intrinsic at every relevant LOB update (taking into account technical delays), with the final schedule determined by the RI.
Leveraging the speed of our proposed DP method, we present full-year RI trading results under a standard battery and simulation setup for the year 2021 in Figure 3. This figure compares rewards, of RI settings outlined in Section 4.1 with a value function discretization setting of $m=11$ and a technical delay of $200$ ms plus measured actual solve times. We evaluate performance starting from the baseline with a solve-frequency of $60$ minutes, i.e. solving the intrinsic optimization once every hour, resulting in a reward of €$221$k (8.7k orders submitted at the exchange) and a simulation runtime of $5.6$ minutes, down to solving it with every relevant order book update (apart from the waiting durations introduced previously) with a reward of €$349$k (30k orders submitted at the exchange) and a simulation runtime $86$ minutes. We define any order message submitted to the LOB as relevant, if it is placed at the head of its corresponding bid/ask stack. These results clearly demonstrate that solving the intrinsic at higher frequencies leads to progressively increasing profits as the trading frequency rises. In this case, a $58\%$ increase between our slowest and quickest solve frequency. We observe a clear increase in the slope of the cumulative reward starting in October 2021. This originates most likely from higher and more volatile gas prices in Europe during this time, which subsequently translates to more extreme intraday electricity prices, which increases the revenue potential of storage.
Figures 4 and 5 illustrate RI decisions, using the same settings and our quickest strategy, solving the intrinsic at every relevant LOB update. Figure 4 provides an overview of the annual battery state-of-charge (SoC) schedule shaped by the RI’s trading decisions, along with the daily trading volumes. We observe a distinct average SoC pattern, suggesting a strong daily periodicity of arbitrage opportunities, with higher charge levels during the early morning and afternoon hours. This can be seen by looking at the trading decisions in the central heat map of trading decisions as well as the mean SoC profile.
The lower plot shows the frequency of different SoC levels throughout the year. Given that our storage has duration of one hour, one would expect the storage to be either completely full or completely empty most of the time. However, the SoC occupies levels in between these extremes surprisingly often. This is most likely caused mainly by complex trading trading behavior induced by nonlinearies in order book-based trading, which makes simple bang-bang strategies suboptimal. This highlights the difference between a naive price taking implementation of the rolling intrinsic and our more realistic approach.
Lastly, when looking at the daily traded energy over the year on the top of the figure, we see an increased trading activity towards the end of the year, which fits the increased slope of cumulative profits for this time period observed in Figure 3.
Figure 4: Yearly heatmap of the final battery schedule for the year 2021 determined by the RI. The upper plot shows the final state of charge (SoC) of the battery for each hour of the year, with the color gradient defined in the lower plot. The gray areas show the total traded energy per day and the yearly mean SoC profile, while the dots in the upper inset represent the largest trade per day submitted by the RI. We observe a clear preference for charging the battery in the early morning and afternoon hours. The lower plot shows the logarithmic distribution of SoC states of the battery, as determined by the RI’s schedule. Figure adapted from (Brudermueller and Kreft, 2023).
Figure 5 shows the number of orders executed by the RI, classified by their proximity to physical delivery. Comparing this to Figure 2 showing the number of orders placed in the market, we see that the pattern is much more uniform. This is likely due to the fact that the rolling intrinsic tries to capitalize on time spreads which often induces trading pairs of products with large differences in the time to delivery, e.g., when buying cheap in the early morning to sell expensive in the evening. This induces the RI to frequently trade products that are not yet liquid, which is one of the biggest weaknesses induced by the myopic nature of the policy.
Figure 5: Order submission-time analysis of the RI simulated over the entire year 2021.
4.4 Parameter Evaluation
The DP rolling intrinsic optimization relies on numerous parameters, each influencing trading behavior and solution accuracy in different ways. This section compares key storage and optimization parameters, demonstrating that our year-long trading results remain robust and that the parameters’ effects align with expectations.
To achieve this, we compare the effects of varying injection-efficiencies (assuming $\eta^{+}=\eta^{-}$), linear degradation costs $\nu_{\text{deg}}$ and maximum storage capacities $\bar{s}$, each for three different equidistant DP value function grids $G$ with granularity $|G|=m$. Table 3 provides a shortened overview of yearly rewards for each of the optimizations. Two clear trends can be observed as expected: A finer grid $G$ (larger $m$) results in more profitable trading as the solution is more accurate, and more penalizing injection efficiencies $\eta$ and linear degradation costs $\nu$ result in lower rewards because fewer trading opportunities become profitable. Both parameters have a significant effect on profits, with the most favorable settings nearly doubling the profits of the least favorable setting. This is especially interesting for the case of degradation cost, which are likely to decrease significantly in the near future, thereby greatly increasing the profitability of short-term trading on intraday markets.
Additionally, batteries with a smaller maximum storage capacity produce larger profits per MW storage capacity, as the increased trading flexibility of larger batteries profits from the same most profitable trading opportunities as with a smaller battery, and only adds less profitable trades on top. For similar reasons, a battery with a duration of $2$h yields lower profits, but more than half of the profits observed for our standard setup with a duration of $1$h.
$\eta^{+}=\eta^{-}$
Reward [€] 0.9 0.95 0.99 1.0
DP-11 277000 349000 441000 478000
DP-51 278000 351000 444000 482000
DP-101 279000 352000 442000 483000
$\nu_{\textbf{deg}}$ [€/MWh]
Reward [€] 8 4 2 0
DP-11 263000 349000 416000 434000
DP-51 264000 351000 419000 439000
DP-101 265000 352000 419000 441000
$\bar{s}$ [MWh]  1h-battery
Reward/$\bar{s}$ 5 10 20 40
DP-11 36600 34900 32600 29700
DP-51 36900 35100 33200 30400
DP-101 36900 35200 33400 30600
$\bar{s}$ [MWh]  2h-battery
Reward/$\bar{s}$ 5 10 20 40
DP-11 27300 27000 25700 23900
DP-51 27100 27400 26200 24600
DP-101 27700 27300 26400 24700
Table 3: A reward comparison of simulation results for the full year 2021, using four different storage parameter settings. All simulation base-settings are equal to the setting defined in 4.1, setting the ping-delay to 200 ms and using actual measured solve-times at execution. Varying storage parameters has the expected effect on yearly rewards, where a finer DP storage discretization generally leads to marginally higher rewards, increasing degradation costs and losses reduces the rewards, and operating a larger storage yields lower rewards per storage capacity. All results are rounded to significance, and the specific battery reward on the right table is given in units of €/MWh.
4.5 Parametrization of the Rolling Intrinsic
The rolling intrinsic strategy is suboptimal due to its myopic nature. In particular, the lack of foresight leads to impatient trading decisions which use flexibility early on, thereby foregoing parts of the storage’s revenue potential. This happens in particular when the RI trades products with limited liquidity, typically long before gate closure, and for rather disadvantageous prices.
This opens the door for parametric variations of the rolling intrinsic policy that aim to improve on these issues by adjusting the trading behavior. However, finding optimal parameters requires extensive tuning efforts and, therefore, only works in combination with an efficient implementation of the policy. Our method therefore opens the door to optimized variations of the rolling intrinsic, which can be trained within a realistic timeframe.
To illustrate this point, we introduce an empirically motivated unitless linear parameter $\phi$ to the intrinsic optimization. It acts as a penalty for products with a large bid-ask spread. We then adapt the optimization (21) by extending $\pi_{t}$ to $\hat{\pi}_{t}=\pi_{t}-(\phi\cdot\delta_{t^{*},t})\mid f_{t}^{0}-f_{t}\mid$, with $\delta_{t^{*},t}$ defined as the big-ask spread of product $t$ at time $t^{*}$. Intuitively, it becomes clear that $\phi$ incentivizes the rolling intrinsic to trade at times of higher market liquidity, i.e. at times closer to delivery, to reduce the chance of trading less profitably too early.
The increased speed of our DP solution method allows us to quickly search for optimal parameters $\phi$. For the sake of simplicity and just to show a proof of concept, we fix the parameters $\phi$ at constant values for each trading month, and use the previous month $m-1$ to train $\phi$ for a given month $m$. When evaluating this strategy over the entire year 2021, this sliding window parameter training results in $12$ separate training windows. We train the parameters on the highest intrinsic-solve frequency (see section 4.3) using DP and simulation settings outlined in Section 4.1, which represents the most realistic training scenario. We treat our monthly simulation rewards as a black box function $f(\phi)$ and, given its simple dependence on a singular parameter, use Brent’s method to find its maximum.
The resulting optimal parameters for the year 2021 have an average penalty of $\bar{\phi}=2.1\pm 1.6$. Comparing year-long out-of-sample results of the newly trained policy yields an increase of $8.4$% in trading reward over the entire year 2021, accumulating €$376$k, compared to the €$347$k reward of the standard policy with the penalty set to 00, i.e., $\phi\equiv 0$.
5 Conclusion
In this paper, we present a novel method to compute the rolling intrinsic policy for trading with a battery on continuous intraday markets. Our method achieves a solution speed improvement of up to three orders of magnitude while maintaining a sufficient level of accuracy. Our work lays the foundation for extensive future research through the extension and adaptation of our rolling intrinsic formulation and implementation, both for researchers and industry.
Our findings demonstrate the importance of strategies that take into account the exact workings of the continuous intraday market and at the same time can be executed fast enough to keep up with the millisecond pace of the market. We show in a realistic backtesting that a solution frequency of at least every second increases rolling intrinsic profits substantially compared to solving the intrinsic only every few minutes or even every hour.
Furthermore, our efficient implementation enables us to backtest the profitability of the rolling intrinsic policy over extended time periods, which requires a substantial number of intrinsic solves across the simulation period, making traditional MILP or LP methods impractical due to their comparatively slow solution time. This also makes it possible to train parametric extensions of the rolling intrinsic that correct some of its shortcomings, as we demonstrate by optimizing a penalty for trading illiquid products, which substantially increases trading profits.
An important future extension of our work lies in more sophisticated parametrizations of the rolling intrinsic. Our naive method could easily be adapted to multiple parameters, affecting various aspects of the RI’s trading behavior, and utilizing the speed of our optimization method. Furthermore, future work could naturally extend the trading to a multi-market scenario, where additional decision variables expand the scope of optimization to multiple subsequent electricity markets, e.g. the reserve or day-ahead markets. Finally, extending our intrinsic optimization with dynamic price forecasts would greatly improve the RI’s profitability, as the optimization would be able to anticipate future price changes. Fine-tuning forecasts to BESS applications would then be a logical utilization of the speed of our method.
Appendix A Code Publication
We publish our method as a Python package BitePy (Battery Intraday Trading Engine), where users can preprocess their intraday market data, set battery and DP parameters, run simulations and analyze results. It is hosted on PyPi and can be easily installed via pip install bitepy. Detailed documentation and tutorials on the package can be found on GitHub.
Appendix B Acknowledgements
The authors thank Markus Kreft and Patrick Langer for the helpful methodological discussions and for support with some of the figures. In addition, we thank Simon Hirsch for the factual checks, general feedback and for the support in wrapping the Package to Python and setting up the documentation.

### v3

@@
 A detailed backtest over a full year of German order book data demonstrates that the proposed dynamic programming formulation does not reduce trading profits and enables the policy to react to every relevant order book update, enabling realistic rapid backtesting.
-Our results show the significant revenue potential of high-frequency trading: our policy earns $58$% more than when re-optimizing only once every hour and $14$% more than when re-optimizing once per minute, highlighting that profits critically depend on trading speed.
-Furthermore, we leverage the speed of our algorithm to train a parametric extension of the rolling intrinsic, increasing yearly revenue by $8.4$% out of sample.
+Our results show the significant revenue potential of high-frequency trading: our policy earns 5858% more than when re-optimizing only once every hour and 1414% more than when re-optimizing once per minute, highlighting that profits critically depend on trading speed.
+Furthermore, we leverage the speed of our algorithm to train a parametric extension of the rolling intrinsic, increasing yearly revenue by 8.48.4% out of sample.
 keywords:
 OR in energy , asset-backed trading , short-term power markets , dynamic programming , limit order book
-journal: EJOR
 […]
@@
   4. 4.
-    We parametrize the standard formulation of the RI, and show that our simulation speed can be used to easily find optimal parameters, generating additional profits of $8.5$, by slightly nudging the RI’s trading behavior at no additional risk.
+    We parametrize the standard formulation of the RI, and show that our simulation speed can be used to easily find optimal parameters, generating additional profits of $8.5$ %, by slightly nudging the RI’s trading behavior at no additional risk.
   5. 5.
@@
 2 The European Power Market
-The European electricity markets operate primarily within two major trading regions, each with their own dedicated exchange: the NordPool exchange, covering the Scandinavian region, and the EPEX exchange, which serves most of the rest of Europe.
+The European short-term electricity markets are served by two primary competing power exchanges, Nord Pool and EPEX SPOT.
+While Nord Pool originated in the Nordic region, both exchanges now operate across wide areas of Europe, including the Nordic and Central Western European (CWE) regions.
 The spot markets, which include day-ahead and intraday markets, are key components of the European Single Day-Ahead Coupling (SDAC) and Single Intraday Coupling (SIDC) initiatives.
@@
 Figure 1 visualizes the process of a typical submission of a new limit order to the limit order book (LOB) of a product, i.e., delivery hour.
+[…]
 Figure 1: An exemplary state of the LOB at $T=1$ presented in the left panel.
-The right panel depicts the clearing of a new hypothetical buy order: A new buy order with a price of 50 €, which is higher than the lowest ask price, is added to the book at $T=2$.
-The quantity of this buy order (30 MW) is then cleared against the cheapest possible offers until either the whole order is fulfilled (as is the case in the figure) or until there are no offers with lower prices left.
-In this example, $22$ MW out of $30$ MW are cleared against the sell order with price 37 € and the remaining 30 MW $-\ 22$ MW $=8$ MW are cleared against the sell order with price 45 €.
-The remaining quantity of 12 MW $-\ 8$ MW $=4$ MW of the latter order stays in the order book.
+The right panel depicts the clearing of a new hypothetical buy order: A new buy order with a price of 50 €/MWh, which is higher than the lowest ask price, is added to the book at $T=2$.
+The quantity of this buy order (30 MWh) is then cleared against the cheapest possible offers until either the whole order is fulfilled (as is the case in the figure) or until there are no offers with lower prices left.
+In this example, $22$ MWh out of $30$ MWh are cleared against the sell order with price 37 €/MWh  and the remaining 30 MWh 22-\ 22 MWh =8=8 MWh are cleared against the sell order with price 45 €/MWh.
+The remaining quantity of 12 MWh 8-\ 8 MWh =4=4 MW of the latter order stays in the order book.
 Note that the clearing is instantaneous, i.e., columns 2–4 in the right panel are purely illustrative and do not correspond to market states that can be observed by traders.
@@
 In the extant literature, the intrinsic problem is set in a perfectly liquid futures market, trading $T\in\mathds{N}$ futures of a commodity that can be stored and which trade for prices $P_{t}$.
-In order not to overcomplicate notation, we assume here that future contracts are for delivery in consecutive time periods $[0,1),[1,2),\dots,[T-1,T)$ and that there are no overlapping contracts.
+In order not to overcomplicate notation, we assume here that future contracts are for delivery in consecutive time periods $0,1),[1,2),\dots,[T-1,T)$ and that there are no overlapping contracts.
 To fix ideas, one can imagine a battery storage system trading futures for delivery in the $24$ hours of the next day.
@@
 More specifically, for a single product $t$ and orders $i\in\mathcal{O}^{+}_{t}$, $j\in\mathcal{O}^{-}_{t}$, buying a quantity $q$ using order $i$ and immediately re-selling the power $\eta^{-}(\eta^{+}q)$ that remains after accounting for efficiency losses using order $j$ yields a positive profit if
-$-qP_{i}+\eta^{-}(\eta^{+}q)P_{j}>0\Leftrightarrow(\eta^{-}\eta^{+})P_{j}>P_{i}.$
+$qP_{i}+\eta^{-}(\eta^{+}q)P_{j}>0\Leftrightarrow(\eta^{-}\eta^{+})P_{j}>P_{i}.$
 Since $P_{i}>P_{j}$ by definition, this only can happen if both prices are negative.
@@
 $\alpha_{t}\in\{0,1\},\;k_{t}\in\mathds{N},\;\nu\in\mathbb{R}_{\geq 0}$ $\forall t\in\mathcal{T}$
-where the $\alpha_{t}$ are binary variables which are equal to $1$ in case the decision for contract $t$ is to buy and 00 if the decision is to sell.
+where the $\alpha_{t}$ are binary variables which are equal to $1$ in case the decision for contract $t$ is to buy and 0 if the decision is to sell.
 Note that, due to the above discussion, strictly speaking, the binary variables $\alpha_{t}$ are only required for hours $t$ where both $\mathcal{O}_{t}^{+}$ and $\mathcal{O}_{t}^{-}$ contain orders with negative prices.
@@
 However, this perspective allows for a faster solution than solving the monolithic MILP or LP as we will argue below.
-We propose solving the problem by an application of the dynamic programming algorithm, discretizing the actions $f_{t}$ into steps $v\cdot u$ ($v\in\mathds{N}$) apart, where $u$ is the minimal tradable quantity.
+We propose solving the problem by an application of the dynamic programming algorithm, discretizing the actions $f_{t}$ into steps $\kappa\cdot u$ ($\kappa\in\mathds{N}$) apart, where $u$ is the minimal tradable quantity.
 Furthermore, we introduce a function $\pi_{t}$ that encodes the current order book information and returns the cost or revenue of buying or selling on the intraday market, taking into account trading costs.
@@
 However, due to efficiencies $\eta^{\pm}<1$, the trades that live on a finite grid translate to storage levels that do not sit on a regular grid, especially after several rounds of injections and withdrawals.
-Therefore, unlike classic applications of the dynamic programming algorithm, we do not define $V_{T}$ to have a finite domain, but instead on the whole interval of possible storage states $[0,\bar{s}]$.
-In order to obtain the value function, we discretize the possible storage states $[0,\bar{s}]$ to a finite grid $G\subset[0,\bar{s}]$ and $\lvert G\rvert=m$.
+Therefore, unlike classic applications of the dynamic programming algorithm, we do not define $V_{T}$ to have a finite domain, but instead on the whole interval of possible storage states $0,\bar{s}]$.
+In order to obtain the value function, we discretize the possible storage states $0,\bar{s}]$ to a finite grid $G\subset[0,\bar{s}]$ and $\lvert G\rvert=m$.
 We then approximate the value function $V_{T}$ by evaluating it on the grid and linearly interpolating for points $s\notin G$.
@@
 Algorithm 2 DP at runtime $t^{*}$
-Input: $f_{t}^{0}$, $\mathcal{O}_{t}$ for all tradable products $t\in\mathcal{T}$, other parameters $(v,s_{0},\nu,G,\dots)$
+Input: $f_{t}^{0}$, $\mathcal{O}_{t}$ for all tradable products $t\in\mathcal{T}$, other parameters $(\kappa,s_{0},\nu,G,\dots)$
 Output: Updated market positions $f_{t}$
@@
 Proof.
-Suppose that the initial state of charge (SoC) is $s_{0}$, we set $v=1$ and look at the simplest case of $f_{t}^{0}=0$ for all $t\in\mathcal{T}$.
+Suppose that the initial state of charge (SoC) is $s_{0}$, we set $\kappa=1$ and look at the simplest case of $f_{t}^{0}=0$ for all $t\in\mathcal{T}$.
 The possible storage states at the beginning of stage $t_{0}+1$ are
@@
 with
-$\mathcal{K}_{t}^{+}$ =$=\left\{k\in\mathds{N}_{0}:k\leq\frac{(\bar{s}-s_{0})+u\max(\mathcal{K}_{t-1}^{-})(\eta^{-})^{-1}}{u\eta^{+}}\right\}$
-$\mathcal{K}_{t}^{-}$ =$=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\eta^{-}\left(s_{0}+\eta^{+}u\max(\mathcal{K}_{t-1}^{+})\right)\right\}.$
+$\mathcal{K}_{t}^{+}$ $=\left\{k\in\mathds{N}_{0}:k\leq\frac{(\bar{s}-s_{0})+u\max(\mathcal{K}_{t-1}^{-})(\eta^{-})^{-1}}{u\eta^{+}}\right\}$
+$\mathcal{K}_{t}^{-}$ $=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\eta^{-}\left(s_{0}+\eta^{+}u\max(\mathcal{K}_{t-1}^{+})\right)\right\}.$
 Note that the way $G_{t}$ and $\mathcal{K}_{t}^{\pm}$ are defined takes care of repeated charging and discharging in periods $t_{0},\dots,t$ and therefore covers all possible storage states. ∎
@@
 We assume efficiency losses of $\eta^{+}=\eta^{-}=0.95$, i.e., a round trip efficiency of $\eta^{+}\eta^{-}\approx 0.9$ similar to Cole and Karmakar (2023); Regelleistung Online (2024).
-We use linear degradation costs of $\nu_{\text{deg}}=4$/MWh, i.e., a round-trip degradation cost of $8$/MWh for every MWh of electricity sold by the battery.
-We motivate this value by Goldman Sachs (2023) who project the cost of battery replacement in 2030 to be around $60$k €/ MWh, a 10-year battery warranty, and approximately $2$ charging cycles per day.
+We use linear degradation costs of $\nu_{\text{deg}}=4$ €/MWh, i.e., a round-trip degradation cost of $8$ €/MWh for every MWh of electricity sold by the battery.
+We motivate this value by Goldman Sachs (2023) who project the cost of battery replacement in 2030 to be around 6060k €/ MWh, a 10-year battery warranty, and approximately $2$ charging cycles per day.
 Dividing the replacement costs by cycles gives the cost of one full cycle, which allows us to infer the induced per MWh degradation cost.
-Furthermore, we choose $\nu_{\text{trade}}=0.09$/MWh, reflecting the current EPEX volume-based trading fees per each matched order.
-In total, the variable cost $\nu=\nu_{\text{trade}}+\nu_{\text{deg}}$ thus takes the value of $\nu=4.09$/MWh.
+Furthermore, we choose $\nu_{\text{trade}}=0.09$ €/MWh, reflecting the current EPEX volume-based trading fees per each matched order.
+In total, the variable cost $\nu=\nu_{\text{trade}}+\nu_{\text{deg}}$ thus takes the value of $\nu=4.09$ €/MWh.
+This variable cost $\nu$ is always present in the optimization, while $\nu_{\text{trade}}$ is deducted from profits after each transaction.
+The degradation cost $\nu_{\text{deg}}$ is subtracted only once, after each product’s final schedule is set at gate closure.
 All results are generated using historical 2021 order book data from the EPEX Spot continuous intraday market for Germany, focusing on hourly products.
-The exchange allows for price increments of $0.01$ and volume increments at $0.1$ MWh, defining the minimum action size on the market as $u=0.1$ MWh.
-The discretization of the dynamic programming action is defined at the most accurate minimum value of $v=1$ throughout.
+It is not possible ex post to simulate iceberg orders correctly given the EPEX dataset, so we omit them in our simulations.
+The exchange allows for price increments of $0.01$ € and volume increments at $0.1$ MWh, defining the minimum action size on the market as $u=0.1$ MWh.
+The discretization of the dynamic programming action is defined at the most accurate minimum value of $\kappa=1$ throughout.
 Clearly, in real-life trading, there will always be some delay between two consecutive optimizations.
@@
 $t^{*}$$t^{*}+\Delta t$Solve TimeTechnical DelayWaiting Duration
-If the combined delay exceeds $1$ ms, this has two effects on trading: Firstly, there is the risk that an order placed by our policy may not be matched since its intended counterpart either expired or was cleared against other market participants’ orders in the meantime.
+If the combined delay exceeds $1$ ms (the tick duration of the EPEX Spot market), this has two effects on trading: Firstly, there is the risk that an order placed by our policy may not be matched since its intended counterpart either expired or was cleared against other market participants’ orders in the meantime.
 We deal with this by submitting all-or-none orders that expire if they are not immediately cleared, which might lead to non-physical positions, i.e., market positions that cannot be matched by a feasible schedule of injections and withdrawals.
@@
 Finally, we remark that when backtesting, we correctly account for changes in the order book caused by our actions.
-This in particular means that orders cleared against the orders of the policy are available shorter than in the actual history, and consequently that subsequent clearing of orders by other markets participants changes.
+This in particular means that EPEX orders cleared against the orders of our policy are available shorter than in the actual history, and consequently that subsequent clearing of orders by other market participants changes.
+This introduces a strong path-dependency of varying trading policies.
 Although we take great care to model the market correctly, a limitation of the backtesting experiments is that, by the very nature of our analysis and the available data, we cannot take into account the effect that the orders placed by the strategy would have had on the behavior of other market participants.
@@
 Due to the simple nature of each single intrinsic optimization, solving the MILP on multiple threads does not lead to faster simulation times.
+Additionally, to provide a fair comparison, we set our our MILP solver’s time-limit to 2 seconds (MIP gap 10%) and warmstart each optimization with the null-action, i.e.
+the solution of not changing the current battery schedule.
 We publish or code in the Python package BitePy, which allows users to easily run our DP simulations over a Python interface.
@@
 For this reason, although inexact in solving single intrinsic problems, the DP-based rolling intrinsic strategy can be more profitable than the MILP formulation, as evidenced by the February and April results in Table 2.
-We choose the most clinical setting for our comparison, by artificially fixing the solve time and technical delay to 00 ms for both approaches.
+We choose the most clinical setting for our comparison, by artificially fixing the solve time and technical delay to 0 ms for both approaches.
 This results in a policy that ignores technical delays and solves the intrinsic optimization for each relevant order book update in such a way that the resulting all-or-none limit orders are matched instantaneously with the exchange after each solve.
@@
 DP-101 11161 0.43 2 901000 1873
-DP-51 11158 0.22 2 900900 1884
+DP-51 11158 0.22 2 $900$ 1884
 DP-11 11181 0.05 2 901000 1899
@@
 We evaluate performance starting from the baseline with a solve-frequency of $60$ minutes, i.e.
-solving the intrinsic optimization once every hour, resulting in a reward of €$221$k (8.7k orders submitted at the exchange) and a simulation runtime of $5.6$ minutes, down to solving it with every relevant order book update (apart from the waiting durations introduced previously) with a reward of €$349$k (30k orders submitted at the exchange) and a simulation runtime $86$ minutes.
+solving the intrinsic optimization once every hour, resulting in a reward of €221221k (8.7k orders submitted at the exchange) and a simulation runtime of $5.6$ minutes, down to solving it with every relevant order book update (apart from the waiting durations introduced previously) with a reward of €349349k (30k orders submitted at the exchange) and a simulation runtime $86$ minutes.
 We define any order message submitted to the LOB as relevant, if it is placed at the head of its corresponding bid/ask stack.
@@
 Additionally, batteries with a smaller maximum storage capacity produce larger profits per MW storage capacity, as the increased trading flexibility of larger batteries profits from the same most profitable trading opportunities as with a smaller battery, and only adds less profitable trades on top.
-For similar reasons, a battery with a duration of $2$h yields lower profits, but more than half of the profits observed for our standard setup with a duration of $1$h.
+For similar reasons, a battery with a duration of 22h yields lower profits, but more than half of the profits observed for our standard setup with a duration of 11h.
 $\eta^{+}=\eta^{-}$
@@
 $\bar{s}$ [MWh]  1h-battery
-Reward/$\bar{s}$ 5 10 20 40
+[…]
 DP-11 36600 34900 32600 29700
@@
 $\bar{s}$ [MWh]  2h-battery
-Reward/$\bar{s}$ 5 10 20 40
+[…]
 DP-11 27300 27000 25700 23900
@@
 The increased speed of our DP solution method allows us to quickly search for optimal parameters $\phi$.
-For the sake of simplicity and just to show a proof of concept, we fix the parameters $\phi$ at constant values for each trading month, and use the previous month $m-1$ to train $\phi$ for a given month $m$.
+For the sake of simplicity and just to show a proof of concept, we fix the parameters $\phi$ at constant values for each trading month, and use the previous month m1m-1 to train $\phi$ for a given month $m$.
 When evaluating this strategy over the entire year 2021, this sliding window parameter training results in $12$ separate training windows.
@@
 The resulting optimal parameters for the year 2021 have an average penalty of $\bar{\phi}=2.1\pm 1.6$.
-Comparing year-long out-of-sample results of the newly trained policy yields an increase of $8.4$% in trading reward over the entire year 2021, accumulating €$376$k, compared to the €$347$k reward of the standard policy with the penalty set to 00, i.e., $\phi\equiv 0$.
+Comparing year-long out-of-sample results of the newly trained policy yields an increase of 8.48.4% in trading reward over the entire year 2021, accumulating €376376k, compared to the €347347k reward of the standard policy with the penalty set to 0, i.e., $\phi\equiv 0$.
 5 Conclusion
@@
 An important future extension of our work lies in more sophisticated parametrizations of the rolling intrinsic.
-Our naive method could easily be adapted to multiple parameters, affecting various aspects of the RI’s trading behavior, and utilizing the speed of our optimization method.
+Our naive method could easily be adapted to multiple additional parameters like trading delays, affecting various aspects of the RI’s trading behavior, further utilizing the speed of our optimization method.
 Furthermore, future work could naturally extend the trading to a multi-market scenario, where additional decision variables expand the scope of optimization to multiple subsequent electricity markets, e.g.
@@
 Fine-tuning forecasts to BESS applications would then be a logical utilization of the speed of our method.
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the authors used Gemini 2.5 and ChatGPT in order to improve language and readability.
+After using this tool/service, the author reviewed and edited the content as needed and take full responsibility for the content of the published article.
 Appendix A Code Publication